5 times the sum of a number and -1 is the same as 6 times the number.
Answer by nerdybill(7384)

5 times the sum of a number and -1 is the same as 6 times the number.
.
Let x = the number
then
5(x+(-1)) = 6x
5(x-1) = 6x
5x-5 = 6x
-5 = x
.
Answer: the "number" is -5